OpenStudy (anonymous):

write a rule to describe the transformation thT is a reflection across the y-axis

OpenStudy (kenljw):

y = f(x) f(-x) = f(x) Even function f(-x) = -f(x) Odd function This shows the relation about x = 0, the y-axis